# Instant Runoff Voting and the Reinforcement Paradox

### v2

@@ 3. Necessary and Sufficient Conditions
(Complete Ballots). An election demonstrates a runner-up reinforcement paradox if and only if
@@
    $Y>X-Z_{2}\text{ and }Y>-X_{1}+X_{2}+Z_{1}-Z_{2}.$
If the election satisfies (1) then the runner-up can win a majority in one of the sub-elections and if the election satisfies (2) then the runner-up does not win a majority in either sub-election.
$Z>X-Y_{2},\ Z>X_{1}-X_{2}+Y_{1}-Y_{2},\ Z>-X_{2}+Y,\text{ and }\ Z>\frac{1}{3}(X+Y).$
Proposition 2.
(Partial Ballots). An election demonstrates a runner-up reinforcement paradox if and only if
@@
    $Y^{\prime}>X^{\prime}\text{ and }Y^{\prime}>-X_{1}+X_{2}+Z^{\prime},$ or
@@
    $Y^{\prime}>X^{\prime}-Z_{2},Y^{\prime}>-X_{1}+X_{2}+Z_{1}-Z_{2}$, and $Y^{\prime}>\frac{1}{3}(-X_{1}+2X_{2}+2X_{3}+Z_{1}-Z_{2}).$
If the election satisfies (1) then runner-up can win a majority in one of the sub-elections and if the election satisfies (2) then the runner-up does not win a majority in either sub-election.
$Z^{\prime}>X^{\prime}-Y_{2},\ Z^{\prime}>X_{1}-X_{2}+Y_{1}-Y_{2},\ Z^{\prime}>-X_{2}+Y^{\prime},\ Z^{\prime}>\frac{1}{3}(X^{\prime}+Y^{\prime}).$
The conditions in Proposition 2 reduce to those in Proposition 1 when $X_{3}=Y_{3}=Z_{3}=0$ for both runner-up and plurality loser reinforcement paradoxes. The third condition $Y^{\prime}>\frac{1}{3}(-X_{1}+2X_{2}+2X_{3}+Z_{1}-Z_{2})$ in Case (2) of Proposition 2 reduces to $Y>\frac{1}{3}(-X_{1}+2X_{2}+Z_{1}-Z_{2})$ and does not have an analogue in Proposition 1 because it is implied by the other conditions in Case (2). To see this, note that
$X_{1}-2X_{2}$ $+3Y-Z_{1}+Z_{2}=(X-Y+Z_{1}-Z_{2})$
$+2(-X+Y+Z_{2})+(X_{1}-X_{2}+Y-Z_{1}+Z_{2})+X_{1}+(Y-Z).$
But $-X+Y+Z_{2}>0$ and $X_{1}-X_{2}+Y-Z_{1}+Z_{2}>0$ by the remaining two conditions in Case (2) and $Y-X>0$ by assumption. Hence $X_{1}-2X_{2}+3Y-Z_{1}+Z_{2}>0$, which implies $Y>\frac{1}{3}(-X_{1}+2X_{2}+Z_{1}-Z_{2})$.
Before sketching the proofs of the propositions, we make a few observations. Firstly, a runner-up and plurality loser reinforcement paradoxes can occur when the IRV and plurality winners are distinct or when they coincide, assuming either complete or partial ballots. Secondly, in Case (2) of the runner-up paradox and in the plurality loser paradox it is easily seen that the runner-up (plurality loser respectively) wins both sub-elections only when different candidates are eliminated in the two sub-elections.
Thirdly, in Case (1) of the runner-up paradox, the runner-up must be the plurality winner, and the proof of the propositions show that if the two conditions are met, the runner-up wins a majority of first-place votes in one sub-election and wins in two rounds with the IRV winner being the plurality loser in the other sub-election. Note in Case (1), if $X_{1}>X_{2}$ then $Y^{\prime}>-X_{1}+X_{2}+Z^{\prime}$ is automatically satisfied. Thus, if among the voters whose first choice is the IRV winner, more voters second-rank the runner-up than second-rank the plurality loser, the paradox occurs whenever the runner up is the plurality winner ($Y^{\prime}>X^{\prime}$ or $Y>X$ in the complete ballot case).
We sketch the proof of Proposition 1 for the plurality loser. The remaining proofs are similar.
Proof.
Let $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$ be a partition of the original election into two sets of ballots in which $C$ is the IRV winner. Let $\mathcal{B}_{1}$ have voter profile as shown in Table 3 where $0\leq x_{1}\leq X_{1}$ replaces $X_{1}$, $0\leq x_{2}\leq X_{2}$ replaces $X_{2}$ and so on.
Candidate $C$ cannot have a majority of first-place votes in both sub-elections (otherwise it would have a majority in the original election). We claim that $C$ cannot have a majority of first-place votes in either sub-election. To see this, suppose $C$ has a majority in $\mathcal{B}_{1}$. If $A$ is the plurality loser in $\mathcal{B}_{2}$, then
$x_{1}+x_{2}<z_{1}+z_{2}-y_{1}-y_{2}\mbox{ and }X_{1}+X_{2}-x_{1}-x_{2}<Z_{1}+Z_{1}-z_{1}-z_{2}.$
Adding these inequalities yields $X_{1}+X_{2}<Z_{1}+Z_{2}-y_{1}-y_{2}<Z_{1}-Z_{2}$, which is not the case. The same contradiction arises if $B$ is the plurality loser in $\mathcal{B}_{2}.$ This proves the claim.
Thus, $C$ must win in the second round after the elimination of a plurality loser in both sub-elections. The two sub-elections cannot have the same plurality loser (otherwise that candidate would be a plurality loser when the ballots were combined). So assume $B$ is the plurality loser in $\mathcal{B}_{1}$ and $A$ is the plurality loser in $\mathcal{B}_{2}.$ This yields the following set of inequalities.
(1) $y_{1}+y_{2}$ $<x_{1}+x_{2}$
(2) $y_{1}+y_{2}$ $<z_{1}+z_{2}$
(3) $x_{1}+x_{2}+y_{1}$ $<z_{1}+z_{2}+y_{2}$
(4) $X_{1}+X_{2}-x_{1}-x_{2}$ $<Y_{1}+Y_{2}-y_{1}-y_{2}$
(5) $X_{1}+X_{2}-x_{1}-x_{2}$ $<Z_{1}+Z_{2}-z_{1}-z_{2}$
(6) $Y_{1}+Y_{2}-y_{1}-y_{2}+X_{1}-x_{1}$ $<Z_{1}+Z_{2}-z_{1}-z_{2}+X_{2}-x_{2}$
Note that since $Z_{1}+Z_{2}<Y_{1}+Y_{2}$, Equations (1) and (5) imply
$X_{1}+X_{2}-x_{1}-x_{2}<Z_{1}+Z_{2}-z_{1}-z_{2}<Y_{1}-Y_{2}-y_{1}-y_{2}.$
So we can eliminate (4). Recalling that $z=z_{1}+z_{2}$ and $Z=z_{1}+z_{2},$ and incorporating the conditions $0\leq y_{1}\leq Y_{1}$, etc. into the constraints,we can rewrite these inequalities as
$\max\{0,$ $X_{1}-X_{2}+Y_{1}+Y_{2}-Z-x_{1}+x_{2}-y_{2}+z\}<y_{1}<\min\{Y_{1},$
$x_{1}+x_{2}-y_{2},z-y_{2},-x_{1}-x_{2}+z+y_{2}\}$
and
$0<z<\min\{Z,-X_{1}-X_{2}+Z+x_{1}+x_{2}\}$
Since every lower bound of $y_{1}$ must be bounded above by every upper bound of $y_{1},$ these bounds on $y_{1}$ lead to the following conditions
00 0 $<x_{1}+x_{2}-y_{2}$
$\quad\Rightarrow y_{2}<x_{1}+x_{2}$
00 0 $<z-y_{2}$
$\quad\Rightarrow y_{2}<z$
00 0 $<-x_{1}-x_{2}+z+y_{2}$
$\quad\Rightarrow x_{1}+x_{2}-y_{2}<z$
$X_{1}-X_{2}+Y_{1}+Y_{2}-Z-x_{1}+x_{2}-y_{2}+z$ $<Y_{1}$
$\quad\Rightarrow z<-X_{1}+X_{2}-Y_{2}+Z+x_{1}-x_{2}+y_{2}$
$X_{1}-X_{2}+Y_{1}+Y_{2}-Z-x_{1}+x_{2}-y_{2}+z$ $<x_{1}+x_{2}-y_{2}$
$\quad\Rightarrow z<-X_{1}+X_{2}-Y_{1}-Y_{2}+Z+2x_{1}$
$X_{1}-X_{2}+Y_{1}+Y_{2}-Z-x_{1}+x_{2}-y_{2}+z$ $<z-y_{2}$
$\quad\Rightarrow X_{1}-X_{2}+Y_{1}+Y_{2}-Z+x_{2}<x_{1}$
$X_{1}-X_{2}+Y_{1}+Y_{2}-Z-x_{1}+x_{2}-y_{2}+z$ $<-x_{1}-x_{2}+z+y_{2}$
$\quad\Rightarrow\frac{1}{2}(X_{1}-X_{2}+Y_{1}+Y_{2}-Z)+x_{2}<y_{2}$
Adding the new conditions on $z$ into the bounds on $z$, we have
$\max\{0,\,$ $y_{2},\,x_{1}+x_{2}-y_{2}\}<z<\min\{Z,-X_{1}-X_{2}+Z+x_{1}+x_{2},$
$-X_{1}+X_{2}-Y_{2}+Z+x_{1}-x_{2}+y_{2},-X_{1}+X_{2}-Y_{1}-Y_{2}+Z+2x_{1},\}$
These conditions, along with the other conditions above in which $z$ is eliminated, only involve $z,y_{2},x_{1},x_{2}$ and the parameters of the original equation $X_{1},X_{2}$, etc. Requiring each lower bound of $z$ to be bounded above by every upper bound of $z$ as we did for $y_{1}$, creates another set of inequalities involving $y_{2},x_{1},x_{2}$ and the parameters $X_{1},X_{2}$, etc. We continue in this way, eliminating the variables $y_{2},x_{1}$ and $x_{2}$ in turn until the remaining conditions involve only $X_{1},X_{2}$, etc. Simplified, these reduce to the inequalities in Proposition 1. ∎
The inequalities created in each step of the iterative process in the proofs of the propositions can be used in reverse to identify election partitions leading to the paradoxes. For instance, from the proof of Case (1) of Proposition 1, we can derive the following bounds for the sub-elections.
  • $\max\{X_{2}+X_{3}-Y^{\prime}+Z^{\prime},\frac{1}{2}(X^{\prime}-Y^{\prime}+Z^{\prime})\}<y^{\prime}<\min\{-X^{\prime}+Y^{\prime}+Z^{\prime}\}$
  • $\max\{X^{\prime}-Y^{\prime}+Z^{\prime},\frac{1}{2}(X^{\prime}-Y^{\prime}+Z^{\prime})+\frac{1}{2}y^{\prime},-Y^{\prime}+Z^{\prime}+y^{\prime}\}<z^{\prime}<\min\{Z^{\prime},X_{1}+y^{\prime},2y^{\prime}\}$
  • $\max\{0,X_{2}-Y^{\prime}+Z^{\prime}+y^{\prime}-z^{\prime}\}<x_{2}<\min\{X_{2},z^{\prime},X_{1}+y^{\prime}-z^{\prime},y^{\prime}-\frac{1}{2}z^{\prime},\frac{1}{2}y^{\prime}\}$
  • $\max\{0,-y^{\prime}+z^{\prime}+x_{2},X_{1}+X_{2}-Y^{\prime}+Z^{\prime}-x_{2}+y^{\prime}-z^{\prime}\}<x_{1}<\min\{X_{1},y^{\prime}-x_{2},z^{\prime}-x_{2}\}$
  • $\max\{0,X^{\prime}-Y^{\prime}+Z^{\prime}-x_{1}-x_{2}+y^{\prime}-z^{\prime}\}<x_{3}<\min\{X_{3},y^{\prime}-x_{1}-x_{2},z^{\prime}-x_{1}-x_{2}\}$
Care must be taken when applying these bounds: the inequalities hold only in the limit as no allowance is made for the discrete mature of ballot counts and the possibility of ties. However, if the $X_{1},X_{2},$ etc. values in the original election are sufficiently distinct, these bounds can be used to identify two sub-elections.
For instance, in the complete ballot case, if $X_{1}=100,X_{2}=105,Y=215,Z_{1}=110$ and $Z_{2}=90$, these conditions allow us to select (among other possibilities)
$x_{1}=99,x_{2}=1,y=106\mbox{ and }z=197,$
which leads to
$X_{1}-x_{1}=1,X_{2}-x_{2}=104,Y-y=109,Z-z=3.$
@@
It is easily seen that $A$ wins the original election, but in $\mathcal{B}_{1}$, candidate $B$ wins in two rounds after $A$ is eliminated and in $\mathcal{B}_{2}$, candidate $B$ has a majority.

### v3

@@ 2. Necessary and Sufficient Conditions
(Complete Ballots). An election demonstrates a runner-up reinforcement paradox if and only if the election satisfies
(1) $Y>X-Z_{2}+1\text{ and }Y>-X_{1}+X_{2}+Z_{1}-Z_{2}+1.$
That is, conditions (1) are necessary and sufficient for there to exist a partition of the election into two sub-elections in both of which the runner-up is the winner. If in addition,
(2) $Y>X+1\text{ and }Y>-X_{1}+X_{2}+Z+1,$
then there is a partition in which the runner-up wins a majority of first-place votes in one of the sub-elections and wins in the second round in the other sub-election.
(3) $Z$ $>X-Y_{2}+1,$
$Z$ $>X_{1}-X_{2}+Y_{1}-Y_{2}+1,$
$Z$ $>-X_{2}+Y+1,\text{ and }$
$Z$ $>\tfrac{1}{3}(X+Y)+1.$
In this case, there is a partition in which the plurality loser wins in the second round in both sub-elections.
Proposition 2.
(Partial Ballots). An election demonstrates a runner-up reinforcement paradox if and only if the election satisfies
(4) $Y^{\prime}$ $>X^{\prime}-Z_{2}+1,$
$Y^{\prime}$ $>-X_{1}+X_{2}+Z_{1}-Z_{2}+1,\text{ and}$
$Y^{\prime}$ $>\tfrac{1}{3}(-X_{1}+2X_{2}+2X_{3}+Z_{1}-Z_{2})+1.$
That is, conditions (4) are necessary and sufficient for there to exist a partition of the election into two sub-elections in both of which the runner-up is the winner. If in addition,
(5) $Y^{\prime}>X^{\prime}+1\text{ and }Y^{\prime}>-X_{1}+X_{2}+Z^{\prime}+1,$
then there is a partition in which the runner-up wins a majority of first-place votes in one of the sub-elections and wins in the second round in the other sub-election.
(6) $Z^{\prime}$ $>X^{\prime}-Y_{2}+1,$
$Z^{\prime}$ $>X_{1}-X_{2}+Y_{1}-Y_{2}+1,$
$Z^{\prime}$ $>-X_{2}+Y^{\prime}+1,\text{ and}$
$Z^{\prime}$ $>\tfrac{1}{3}(X^{\prime}+Y^{\prime})+1.$
In this case, there is a partition in which the plurality loser wins in the second round in both sub-elections.
In subsequent sections we sometimes consider a “limiting case” where the number of voters approaches infinity. In this case, the values at the top of a preference profile can be considered proportions rather than vote counts, in which case the “+1” term in the above propositions disappears. That is, the above propositions still hold for the limiting case by eliminating the plus 1 term, which corresponds to dividing 1 by a larger and larger number of voters as we take the limit.
The conditions in Proposition 2 reduce to those in Proposition 1 when $X_{3}=Y_{3}=Z_{3}=0$. The condition $Y^{\prime}>\frac{1}{3}(-X_{1}+2X_{2}+2X_{3}+Z_{1}-Z_{2})+1$ in Proposition 2 reduces to $Y>\frac{1}{3}(-X_{1}+2X_{2}+Z_{1}-Z_{2})+1$ and does not have an analogue in Proposition 1 because it is implied by the other conditions. To see this, note that
$X_{1}-2X_{2}$ $+3Y-Z_{1}+Z_{2}-3=(X-Y+Z_{1}-Z_{2})$
$+2(-X+Y+Z_{2}-1)+(X_{1}-X_{2}+Y-Z_{1}+Z_{2}-1)+X_{1}+(Y-Z).$
But $X+Y+Z_{2}-1\geq 0$ and $X_{1}-X_{2}+Y-Z_{1}+Z_{2}-1\geq 0$ by the remaining two conditions in (1), and YX>0Y-X>0 and $X-Y+Z_{1}-Z_{2}>0$ by assumption. Hence $X_{1}-2X_{2}+3Y-Z_{1}+Z_{2}-3>0$, which implies $Y>\frac{1}{3}(-X_{1}+2X_{2}+Z_{1}-Z_{2})+1$.
Before sketching the proof of the propositions, we make a few remarks. First, an election can be susceptible to either a runner-up or plurality loser reinforcement paradox when the IRV and plurality winners in the original election are distinct or when they coincide, for the complete ballot or partial ballot case.
Second, condition (2) clearly can only be satisfied if the IRV runner-up is the plurality winner. In this case, as the proof will demonstrate, if there is a partition in which the runner-up wins a majority of first-place votes in one sub-election, they must win in two rounds in the other sub-election with the original IRV winner being eliminated in round 1.
Third, it is easy to show that it is not possible to have a partition in which the plurality loser has a majority of first-place votes in one sub-election and wins in two rounds in the other sub-election. Instead, a plurality loser reinforcement paradox can only occur when the plurality loser wins in two rounds in both sub-elections.
We prove a portion of Proposition 1, as outlined below. The remaining proofs are analogous (and considerably longer).
Proof.
(of Proposition 1) We prove that the inequalities in (2) of Proposition 1 are necessary and sufficient for there to be a partition in which the runner-up wins a majority of first-place votes in one of the sub-elections and wins in the second round in the other sub-election.
Suppose the original election has a voter profile as shown in Table 3. Since candidate $B$ is not eliminated in either sub-election, we can collapse the two columns in Table 3 that rank $B$ first and simply consider the number of number of voters $Y$ who rank $B$ first and the remaining candidates in any order. Now suppose $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$ are a partition of the election into two sets of ballots. Let the voter profile of $\mathcal{B}_{1}$ be as in Table 3 with $x_{1}$ replacing $X_{1}$, $x_{2}$ replacing $X_{2}$, $y$ replacing $Y$, $z_{1}$ replacing $Z_{1}$ and $z_{2}$ replacing $Z_{2}$, for some $0\leq x_{1}\leq X_{1}$, $0\leq x_{2}\leq X_{2}$, $0\leq y\leq Y,$ etc. Let $\mathcal{B}_{2}$ be the same except with $X_{1}-x_{1},$ $X_{2}-x_{2}$, YyY-y, $Z_{1}-z_{1}$ and $Z_{2}-z_{2}$.
Suppose $B$ wins by a majority of first-place votes in $\mathcal{B}_{2}$. Clearly, it also cannot win by a majority in $\mathcal{B}_{1}$ (otherwise $B$ would have a majority of first-place votes in the original election). So $B$ must win in two rounds in $\mathcal{B}_{1}.$ We claim that $A$ must be the plurality loser in $\mathcal{B}_{1}.$
To see this, suppose $C$ is the plurality loser in $\mathcal{B}_{1}.$ After $C$ is eliminated,
$x_{1}+x_{1}+z_{1}<y+z_{2}.$
On the other hand, since $B$ wins in a majority in $\mathcal{B}_{2},$
$(X_{1}-x_{1})+(X_{2}-x_{2})+(Z_{1}-z_{1})+(Z_{2}-z_{2})<Y-y.$
Solving both of these inequalities for $z_{1}$, and recalling that $z_{1}$ is an integer, we have
$X_{1}+X_{2}-Y+Z_{1}+Z_{2}-x_{1}-x_{2}+y-z_{2}+1\leq z_{1}\leq-x_{1}-x_{2}+y+z_{2}-1,$
which implies
$X_{1}+X_{2}-Y+Z_{1}+Z_{2}+2\leq 2z_{2}.$
Since $z_{2}\leq Z_{2}$, this implies $X_{1}+X_{2}-Y+Z_{1}+Z_{2}+2\leq 2Z_{2}$ or
$X_{1}+X_{2}+Z_{1}+2\leq Y+Z_{2},$
which is a contradiction, since $A$ is the IRV winner in the original election. This proves the claim.
Thus $A$ must be the plurality loser in $\mathcal{B}_{1}$ with $B$ winning in round 2. Let $z=z_{1}+z_{2}$ be the number of voters in $\mathcal{B}_{1}$ who first-rank $C$ and let $Z-z=(Z_{1}-z_{1})+(Z_{2}-z_{2})$ be the number of analogous voters in $\mathcal{B}_{2}$. We require (recalling that $X=X_{1}+X_{2}$)
(7) $x_{1}+x_{2}$ $<y$
(8) $x_{1}+x_{2}$ $<z$
(9) $x_{2}+z$ $<y+x_{1}$
(10) $(X-x_{1}-x_{2})+(Z-z)$ $<(Y-y).$
We simplify these inequalities in a series of steps in which we sequentially eliminate variables. In the first step, we collect lower and upper bounds for $y$.
Step 1: Solving (7), (8) and (10) for $y$ and adding the boundary constraints $0\leq y\leq Y,$ we have
$\max\{x_{1}+x_{2}+1,-x_{1}+x_{2}+z+1\}\leq y\leq\min\{Y,-X+Y-Z+x_{1}+x_{2}+z-1\}$
Since each lower bound must be bounded above by each upper bound, this implies
$x_{1}+x_{2}+1$ $\leq Y$ $\Rightarrow$ $x_{2}\leq Y-x_{1}-1$
$x_{1}+x_{2}+1$ $\leq-X+Y-Z+x_{1}+x_{2}+z-1$ $\Rightarrow$ $z\geq X-Y+Z+2$
$-x_{1}+x_{2}+z+1$ $\leq Y$ $\Rightarrow$ $z\leq Y+x_{1}-x_{2}-1$
$-x_{1}+x_{2}+z+1$ $\leq-X+Y-Z+x_{1}+x_{2}+z-1$ $\Rightarrow$ $x_{1}\geq\frac{1}{2}(X-Y+Z+2).$
Step 2: Collecting all the bounds (including (8)) on $z$, along with the boundary constraint $0\leq z\leq Z$, we have
$\max\{x_{1}+x_{2}+1,X-Y+Z+2\}\leq z\leq\min\{Z,Y+x_{1}-x_{2}-1\}.$
Again, since each lower bound must be bounded above by each upper bound, we have
$x_{1}+x_{2}+1$ $\leq Z$ $\Rightarrow$ $x_{2}\leq Z-x_{1}-1$
$x_{1}+x_{2}+1$ $\leq Y+x_{1}-x_{2}-1$ $\Rightarrow$ $x_{2}\leq\frac{1}{2}(Y-2)$
$X-Y+Z+2$ $\leq Z$ $\Rightarrow$ $Y\geq X+2$
$X-Y+Z+2$ $\leq Y+x_{1}-x_{2}-1$ $\Rightarrow$ $x_{2}\leq-X+2Y-Z+x_{1}-3$
Step 3: Collecting all the bounds on $x_{2}$, we have
$\max\{0\}\leq x_{2}\leq\min\{X_{2},Y-x_{1}-1,Z-x_{1}-1,\frac{1}{2}(Y-2),-X+2Y-Z+x_{1}-3\}.$
Hence
$0$ $\leq Y-x_{1}-1$ $\Rightarrow$ $x_{1}\leq Y-1$
$0$ $\leq Z-x_{1}-1$ $\Rightarrow$ $x_{1}\leq Z-1$
$0$ $\leq\frac{1}{2}(Y-2)$ $\Rightarrow$ $Y\geq 2$
$0$ $\leq-X+2Y-Z+x_{1}-3$ $\Rightarrow$ $x_{1}\geq X-2Y+Z+3$
Step 4: Finally, collecting all the bounds on $x_{1}$, we have
$\max\{0,\frac{1}{2}(X-Y+Z+2),X-2Y+Z+3\}\leq x_{1}\leq\min\{X_{1},Z-1,\}.$
This implies
$0$ $\leq Z-1$ $\Rightarrow$ $Z\geq 1$
$\frac{1}{2}(X-Y+Z+2)$ $\leq X_{1}$ $\Rightarrow$ $Y\geq-X_{1}+X_{2}+Z+2$
$\frac{1}{2}(X-Y+Z+2)$ $\leq Z-1$ $\Rightarrow$ $Y\geq X-Z+4$
$X-2Y+Z+3$ $\leq X_{1}$ $\Rightarrow$ $Y\geq\frac{1}{2}(X+Z+3)$
$X-2Y+Z+3$ $\leq Z-1$ $\Rightarrow$ $Y\geq\frac{1}{2}(X+4)$
Step 5: Collecting all the inequalities relating $X_{1}$, $X_{2}$, $Y$ and $Z$, these reduce to (assuming $Z\geq 4$):
$Y\geq X+2\quad\text{and}\quad Y\geq-X_{1}+X_{2}+Z+2.$
Since $Y$ is an integer, these conditions are equivalent to the conditions $Y>X+1$ and $Y>-X_{1}+X_{2}+Z+1$ in (2) of Proposition 1. ∎
Propositions 1 and 2 identify the necessary conditions on the original election for a runner-up or plurality loser reinforcement paradox to occur but do not identify any partitions. However, the inequalities that arise at each step of the propositions’ proofs can be used to identify election partitions leading to the paradoxes.
For instance, from the proof above of the necessity of (2) of Proposition 1, we have
@@
    $\max\{x_{1}+x_{2}+1,X-Y+Z+2\}\leq z\leq\min\{Z,Y+x_{1}-x_{2}-1\}$
@@
    $\max\{0\}\leq x_{2}\leq\min\{X_{2},Y-x_{1}-1,Z-x_{1}-1,\frac{1}{2}(Y-2),-X+2Y-Z+x_{1}-3\}$
@@
    $\max\{0,\frac{1}{2}(X-Y+Z+2),X-2Y+Z+3\}\leq x_{1}\leq\min\{X_{1},Z-1,\}.$
Thus, if $X_{1}=50,$ $X_{2}=50$, $Y=102$, $Z_{1}=50$ and $Z_{2}=45,$ we can select (among other possibilities), a partition in which the first sub-election is
$x_{1}=48,x_{2}=40,y=89\mbox{ and }z=95,$
and the second partition is
$X_{1}-x_{1}=2,X_{2}-x_{2}=10,Y-y=13,Z-z=0.$
@@
It is easily seen that $A$ wins the original election, but in $\mathcal{B}_{1}$ candidate $B$ wins in two rounds after $A$ is eliminated and in $\mathcal{B}_{2}$, candidate $B$ has a majority.